### v1

# Improving the prediction of individual engagement in recommendations using cognitive models

Abstract.
For public health programs with limited resources, the ability to predict how behaviors change over time and in response to interventions is crucial for deciding when and to whom interventions should be allocated. Using data from a real-world maternal health program, we demonstrate how a cognitive model based on Instance-Based Learning (IBL) Theory can augment existing purely computational approaches. Our findings show that, compared to general time-series forecasters (e.g., LSTMs), IBL models, which reflect human decision-making processes, better predict the dynamics of individuals’ states. Additionally, IBL provides estimates of the volatility in individuals’ states and their sensitivity to interventions, which can improve the efficiency of training of other time series models.
Cognitive Modeling, Instance-based learning, Time-series forecasting, Limited resource allocation, Maternal health, Engagement
copyright: acmlicensedjournalyear: 2024conference: the 6th International Workshop on Health Recommender Systems co-located with the 18th ACM Conference on Recommender Systems; October 18, 2024; Bari, Italyisbn: ccs: Computing methodologies Modeling and simulationccs: Human-centered computing Collaborative and social computing
1. Introduction
Public health programs play an essential role in improving the health outcomes of individuals and communities, often through education and subsequent behavioral change. Some health programs interact with their intended beneficiaries in a broad and infrequent manner. For example, a campaign about the health risks of smoking may address a general population of smokers through scattered advertisements in the media (Nogueira et al., 2018). Others rely on repeated direct interactions with their intended beneficiaries. For example, maternal health programs that send automated messages about exercise and nutrition to enrolled expectant mothers (Lalan et al., 2024). In this case, it is crucial that mothers remain engaged for the duration of the program or as long as possible to receive the maximum benefit. Unfortunately, many such programs face high levels of disengagement and dropout, which severely limit their effectiveness (Amagai et al., 2022).
To reduce dropout, programs can provide interventions designed to increase beneficiary engagement. For example, program staff can make personalized service calls to beneficiaries at risk of dropping out to address concerns about declining participation. However, programs with limited staff and resources cannot intervene on all their beneficiaries. Instead, they must choose a subset of beneficiaries to get the intervention.
Recent approaches to this challenge have adopted the Restless Multi-Armed Bandit (RMAB) resource allocation framework (Killian et al., 2023). Within this framework, the engagement dynamics of each beneficiary (i.e., transition between discrete engagement states) are modeled as a Markov Decision Process (MDP). A centralized planner decides which beneficiaries to intervene upon at each time point based on the dynamics learned from the MDPs. Importantly, beneficiaries can change their engagement state even without an intervention. The most optimal approach in this framework is the Whittle index (Whittle, 1988), which ranks beneficiaries according to the likelihood of becoming and remaining engaged given an intervention.
A key assumption of this framework is that beneficiaries’ dynamics can be modeled as an MDP. However, it is questionable whether these dynamics are indeed Markovian in reality. Recent work demonstrated that these dynamics are often non-Markovian; that is, a beneficiary’s transitions between states (i.e., engaged and disengaged) do depend on their individualized history, and instead developed the Time-series Arm Ranking Index (TARI) which extends RMABs to allow for non-Markovian dynamics and continuous engagement levels (Danassis et al., 2023). According to TARI, each beneficiary is modeled not as a MDP, but using time-series forecasters such as LSTMs (Hochreiter and Schmidhuber, 1997) or transformers (Vaswani et al., 2017). Using these learned time-series models, TARI computes an index that reflects the relative engagement benefit of intervening on a beneficiary versus not intervening at a given time point, and then ranks beneficiaries according to the computed indices.
Within the context described thus far, we propose the use of computational cognitive models of individual beneficiaries instead of LSTMs. Specifically, we will create a model based on a successful account of individual dynamic decision making, Instance-Based Learning Theory (IBLT) (Gonzalez et al., 2003) to represent each beneficiary’s time series activities. Using an IBL model is expected to have at least two potential advantages over general time-series forecasters (TSFs).
First, TSFs need large amounts of data to fine-tune due to the large number of model parameters. Given the relatively sparse amount of data available per beneficiary, existing approaches typically train a single model using data from all available beneficiaries. For instance, Danassis and colleagues (2023) split each beneficiary’s trajectory into a set of fixed-length vectors where each vector represented the intervention history and engagement levels of seven consecutive time steps. Then an LSTM was trained on these vectors, where the task was to predict the engagement level at the next time step (Danassis et al., 2023). In contrast, an IBL cognitive model can be personalized because it can model and predict individual human behavior with less data (Bugbee and Gonzalez, 2022). This is because IBLT makes strong assumptions about the dynamics of human behavior that are derived from prior research on human cognition and decision making. In particular, IBLT accounts for memory effects (i.e., the more recent a data point is, the more it contributes to the prediction) and similarity effects (i.e., engagement levels that are observed under similar contexts should be similar to each other) that general TSFs do not.
Second, the parameters of IBL cognitive models are grounded in psychological constructs such as memory and attention, and are thus more easily interpreted than the parameters of general LSTMs. This advantage in turn enables cognitive models to provide insight about individual differences in cognition that may affect beneficiaries’ behaviors. We demonstrate how these individual cognitive differences can be used to augment existing approaches that rely primarily on general LSTMs.
1.1. Our Contributions
  1. (1)
    Using real-world data from a maternal health program, we show that cognitive models are more accurate compared to LSTMs in capturing the temporal dynamics of individual behaviors and predicting the behaviors of individual beneficiaries.
  2. (2)
    We showcase that personalized cognitive models could reveal individual characteristics through weights fitting. We illustrate how to cluster individuals using learnt characteristics.
  3. (3)
    We show clustering based on learnt individual characteristics could guide and improve the performance of time series models. Specifically, we show that LSTMs trained in clusters better predict behaviors within those clusters compared to LSTMs trained on random samples of beneficiaries.
1.2. Background
1.2.1. Time-series Arm Ranking Index (TARI)
In the context of multiple potential beneficiaries treated as arms in an RMAB, each arm is considered an independent time series (Danassis et al., 2023). We train a model to predict the next state $s_{t+1}$ based on three pieces of information:
  • A historical record (length h) of the past states and actions of the arm (denoted by $\{(s_{i},a_{i})\}_{i<t}^{h}$).
  • The current state (e.g., engagement level) of the arm $s_{t}$.
  • A potential action $a_{t}$, where 1 represents an intervention and 0 represents no intervention.
This training takes place offline. During test time, the model’s parameters are fixed and we use the trained model with iterated multi-step to generate a long-term forecast of future states $s_{t+1},s_{t+2},...,s_{t+H}$ (Taieb et al., 2012). This forecast is then used to calculate the TARI index, which helps us to decide the relative utility of intervening in an arm.
For each arm, we use the TSF model to estimate two values:
  • Time to disengagement with intervention $u_{n}$: This predicts how long it will take for the arm to become non-engaged if we intervene once at the current timestep (t) and then never intervene again.
  • Time to disengagement without intervention $v_{n}$: This predicts how long it will take for the arm to become non-engaged if we did not intervene at all.
The TARI index for an arm is simply the ratio of these two times. A higher TARI index suggests that intervening now would be more beneficial. Finally, similar to the Whittle index, we choose to act on the $k$ arm with the highest TARI index at each timestep:
$\text{TARI}(n)=\frac{u_{n}}{v_{n}}.$
1.2.2. Instance-Based Learning Theory
Instance-Based Learning Theory (IBLT) is a cognitive theory of dynamic decision making, grounded in human learning and memory mechanisms. It has successfully accounted for human behavior in a variety of contexts, ranging from abstract repeated-choice tasks to more real-world search and choice (Nguyen and Gonzalez, 2022) and continuous control tasks (Gonzalez et al., 2003). In particular, in most of these contexts, human participants are required to make a series of decisions while exposed to changing environments. Given IBLT’s previous success in capturing human decision making, we expect that a model based on IBLT would similarly be able to account for the temporal and intervention-related dynamics of beneficiaries’ engagement levels.
At the core of IBLT is the representation of a decision in the form of an instance. Typically, an instance has three parts: the context, the decision itself, and the associated utility. For example, the darkness of the sky would be the context, the choice to bring an umbrella (or not) would be the decision, and the utility may involve the cost of bringing an umbrella added to whether or not you got wet.
IBLT proposes that the agent makes a prediction of the potential utility of a decision based on similar instances stored in memory. For example, it predicts the potential utility of bringing or not bringing an umbrella in situations with similar sky darkness. Decisions are stored as instances in memory and retrieved when the agent is faced with another decision.To predict the utility of each potential action, the past utilities of similar instances are weighted by their probability of retrieval from memory to generate estimates of the expected utility (i.e., a Blended value). The agent then selects the option that has the highest estimated utility. The utilities of instances in memory are updated using the actual outcome observed for a decision made. Thus, the expected initial utility turns into an experienced utility. The updated instances are then bootstrapped to make future decisions.
The contribution of a retrieved instance’s outcome to the final estimated utility of an option depends on the memory activation of the instance. The activation of an instance ($A_{i}$) reflects how readily it comes to mind, which in turn depends on the frequency, recency, and similarity factors according to the following equation as proposed by the ACT-R cognitive architecture (Anderson and Lebiere, 2014).
(1) $A_{i}=ln(\sum(t-t^{\prime})^{-d})+\mu\sum_{j}({w_{j}(Sim(s_{i,j},s_{t,j})-1})+\sigma\xi$
where $d$ is a decay parameter, $\mu$ is a scaling parameter on the overall impact of similarity to activation, $w_{j}$ is the importance (or weight) of an attribute in the similarity calculations, $Sim$ is a similarity function that returns the degree of similarity for a particular attribute $j$ between instance $i$ and the current context, and $\sigma\xi$ is a scaled noise distribution. The activation of an instance is then converted into a retrieval probability as per the Boltzmann softmax function:
(2) $P_{i}=\frac{e^{A_{i}/\tau}}{\sum_{j}e^{A_{j}/\tau}}$
where $\tau$ is the temperature parameter. The expected utility value of an option $(a)$ is then computed using all retrieved instances weighted by their retrieval probabilities through blending as proposed in (Gonzalez et al., 2003; Gonzalez and Dutt, 2011):
(3) $V(a)=\sum_{i}P_{i}u_{i}$
The agent then chooses the option with the highest blended value $V$.
2. Proposed Approach
Consistent with IBLT, we account for engagement at the individual level such that each beneficiary is modeled by an IBL model. Each IBL model is trained through model tracing; at each timestep, the model is given an instance that represents the beneficiary’s recent engagement history and current engagement level (see Figure 1a). This instance is then stored in the model’s memory, and is subject to the same temporal and similarity mechanisms that govern its influence on future estimates of the beneficiary’s engagement level.
For IBL models to act as time-series predictors, we use a model’s blended value of the current engagement context as its prediction of the beneficiary’s engagement level. Each IBL instance has two attributes that are derived purely from the beneficiary’s engagement history: the number of time steps since the last intervention and the engagement level on the immediately prior time step. The IBL model determines the similarity between the current context and each instance in memory according to these two attributes, and these similarities are factored into an instance’s activation according to Equation  1.
As mentioned above, the similarity of two instances depends not only on the similarity between attribute values, but also on the weight or influence of the attribute ($w$ in Equation  1). Psychologically, an attribute weight represents how relevant the agent considers an attribute to make similarity judgments between two contexts. To capture potential individual differences between mothers, we personalized an IBL model for a specific mother by finding the combination of attribute weight values that resulted in the smallest training prediction error.
Given the predictions of engagement level, we then follow the TARI algorithm and generate a ranked list of mothers according to the predicted relative benefit in the improvement in engagement each mother receives from an intervention versus not receiving one.
3. Simulation Setup
3.1. Data
We tested our approach on data from a maternal healthcare program operated in partnership with a NGO - ARMMAN (ARMMAN, 2024). The main objective of the program is to improve the health outcomes of expectant and early mothers in vulnerable communities through the dissemination of automated messages about maternal and infancy care. To help increase the engagement of mothers enrolled in the program, ARMMAN staff can deliver a limited number of interventions per week in the form of voice calls in which mothers have the opportunity to ask questions and interact with ARMMAN staff.
The dataset was collected by ARMMAN in 2022 and contains the data of 12,000 mothers over a 40-week period. Two data points were recorded for each mother per week: whether or not she received a call from an ARMMAN staff and the amount of time she spent listening to automated health messages. The latter is used as a proxy for engagement level, normalized to the range [0, 1]. Although the dataset also contained the demographic information of each mother, it was found that there was little or no prediction benefit in incorporating demographic variables as additional features for time-series forecasters (Danassis et al., 2023). Thus, we focus solely on the behavioral variables (presence of intervention and level of engagement).
For an IBL model to be able to predict the effect of an intervention on a mother’s engagement level, it needs to be trained on a trajectory that includes at least one intervention. However, of the 12,000 mothers, only 5,400 mothers received at least 1 intervention early on in their program tenure. We also assume that mothers receive an intervention (or the equivalent of an intervention) upon their enrollment in the program. Furthermore, because the amount of time needed to train all IBL models scales up with the number of mothers, we decided to explore our approach with a subset of 210 mothers out of the 5,400 mothers who received at least one intervention in their actual trajectories.
Figure 1. a) Transforming a beneficiary’s trajectory into IBL instances and LSTM sliding windows. Each IBL instance represents the engagement context at a given timestep and consists of three parts; the engagement level on the immediately prior timestep, the number of timesteps since the last intervention, and the engagement level on the current timestep. In contrast, the LSTM data was constructed by segmenting trajectories into sliding windows of length 7. b) Training / testing setup for comparing IBL and LSTM methods. Different colors indicate data from individual beneficiaries. Under the IBL procedure, one IBL model is trained per beneficiary. Each trained IBL model is then used to make predictions for their respective beneficiaries during testing. Under the LSTM procedure, one LSTM is trained using data from all beneficiaries. The trained LSTM model is then used to make predictions for all beneficiaries during testing. c) LSTM training / testing setups under various IBL-clusters-related conditions. Different colors indicate different cluster types. Note that the LSTM in the Random condition is trained on a random subset of one-third of the training data as compared to the other training conditions.
[…]
A line graph of the prediction error of the two model types - IBL and LSTM - over time.
3.2. Simulation Procedure
The first 25 weeks of each mothers’ trajectory were designated as training data. Following the approach of previous studies, to train the LSTM model, the training data was reconstructed into sliding windows of 7 consecutive time steps. Then an LSTM model was trained in the entire training dataset (see Figure  1b for a visual description of the differences between the IBL and LSTM training / testing procedure).
For training of the IBL models, an IBL model was trained per mother. Each IBL model traced a mother’s trajectory through each time step up to week 25. The best-fitting profile of attribute weights per mother was determined by performing a grid search for the combination that minimized the weighted sum of one-step prediction errors (loss). The grid was constructed by taking all possible combinations of parameter values in the range (0, 0) to (5, 5) in intervals of 0.5.
If an IBL model with a specific combination of attribute weights is an adequate model of an individual’s behavioral dynamics, then the model should increase its prediction accuracy as it acquires more instances and becomes more aligned in its history with that of the modeled individual. We account for this assumption in the model fitting procedure by weighting earlier prediction errors (e.g., at weeks 1 and 2) less than later ones (e.g., at weeks 24 and 25) according to the following equations:
(4) $loss=\sum_{t}{q_{t}*SE_{t}}$
(5) $q_{t}=\frac{e^{t/10}}{\sum_{i}{e^{t_{i}/10}}}$
where $SE_{t}$ is the squared prediction error and $q_{t}$ is the normalized weight factor for the timestep $t$.
The trained models were then tasked with generating predictions for the next 14 weeks. For the one-step prediction task, the models iteratively generated predictions for the next time step (i.e., for week 26, then for week 27, then for week 28, etc.). Consistent with the fact that the sampled mothers did not receive additional interventions beyond the first 14 weeks of their tenure, we simulated these predictions in the absence of additional simulated interventions, which then allowed us to compare the models’ engagement predictions with the ground truth data.
We compared the IBL-TARI and LSTM-TARI policies with three other baseline policies: (i) random allocation, where each beneficiary has the same probability of receiving an intervention, (ii) round-robin allocation, where a beneficiary that received an intervention will not receive another until all other beneficiaries have also received an intervention, and (iii) no intervention, where no interventions are delivered. Consistent with ARMMAN’s limited resources, all policies are allowed a 3% budget (except the no intervention policy), which corresponds to allocating interventions to 6 beneficiaries.
In order to simulate the effect of an intervention on a beneficiary’s engagement level when they did not receive one in reality, it was necessary to develop a counterfactual generator. We followed the approach adopted in previous studies and trained a separate LSTM model on trajectories from the entire dataset (approximately 12000 beneficiaries), which is far larger than the subset used for our policy simulations. This counterfactual generator was used only when a mother’s simulated trajectory deviated from her ground-truth trajectory (i.e., received a simulated intervention when she did not receive one in reality or not receiving a simulated intervention when she did receive one in reality).
4. Results
4.1. Predicting next-step engagement levels
We begin by presenting the results directly comparing the performance of the personalized IBL models and the LSTM model on the next-step prediction task. Figure 2 shows the prediction error of the two approaches during each of the 14 weeks in the testing period. Clearly, IBL models consistently achieve lower prediction errors than LSTM. About 10% less errors on average (0.23 on average) compared to the LSTM model (0.32 on average).
Figure 2. Next-step prediction error per testing time step (smaller is better). N = 210. The personalized IBL models consistently achieves lower average prediction error (0.23 across all time steps) as compared to the LSTM model (0.32), which translates to a 29% reduction in error.
[…]
A line graph of the prediction error of the two model types - IBL and LSTM - over time.
4.2. Using predictions to inform intervention allocation
Given this result, we hypothesized that the TARI algorithm would benefit from using IBL models to fill the role of its time-series forecaster. More accurate next-step predictions should lead to a more accurate prediction of the relative benefit of intervening versus not intervening on a mother, and thus allow for a more accurate ranking of mothers. Figure 7 (in Appendix) shows the percentage of mothers who have an engagement level of 0.25 and above during each test time step following the various intervention allocation policies. The LSTM-TARI policy outperforms all other policies, particularly later on in the testing period. The average percentage of engaged mothers across testing timesteps for the various policies are as follows: IBL-TARI ($55.58\%$); Round-robin ($55.65\%$); Random ($54.29\%$); LSTM-TARI ($61.05\%$); Control ($46.22\%$). Contrary to our expectations, the IBL-TARI policy does not outperform the LSTM-TARI policy on this performance metric.
4.3. Elucidating individual differences with IBL parameters
Beyond more accurate predictions of engagement, another advantage of the proposed IBL approach over more general TSFs (as exemplified by the LSTM approach) is the ability to identify individual differences in beneficiaries’ behavioral characteristics. Because each beneficiary is modeled by a separate IBL model, the parameter values of each model provide potential insights into how beneficiaries differ from each other in their engagement and intervention dynamics.
Specifically, the attribute weights of each mother’s IBL model trained on data from the first 25 weeks of her tenure can be interpreted as measures of how reliably her engagement changes in response to an intervention (intervention-sensitivity) and how consistently her engagement changes between time steps (transition-consistency).
Figure 3 shows the distribution of mothers according to their attributes weight profiles. From the visual inspection of the data, there appear to be three types of mothers. We confirmed that the optimal number of clusters is between 3 to 4 clusters using multiple measures of cluster goodness (silhouette: 4, within-sum-of-squares: 4, gap statistic: 3). For interpretability, subsequent analyzes assume that mothers are partitioned into three clusters.
Each cluster was labeled according to its defining characteristics. Intervention-sensitive mothers ($N$ = 62) are those characterized by a high intervention lag weight; given that such a mother is $t$ time steps away from her most recent intervention, it is likely that her engagement level will be similar to her engagement level when she was $t$ time steps away from previous interventions. Transition-consistent mothers ($N$ = 66) are those who are characterized by a high previous engagement weight; Since $x$ was the engagement level of such a mother in the immediately previous time step, it is likely that her current engagement level will be similar to her engagement level in the time step following a previous state when her engagement level was similar to $x$. State-stable mothers ($N$ = 82) are those characterized by relatively low values in both dimensions; given that $x$ was the level of engagement of such a mother in the immediately preceding time step, its current level of engagement is likely to be similar to $x$.
Figure 3. Distribution of attribute weight profiles. Each dot represents the IBL model personalized to a particular beneficiary (mother). Points are slightly jittered to better visualize density. Lighter points indicate training set mothers while darker points indicate testing set mothers. Beneficiaries are clustered into 3 clusters according to a k-means algorithm. Larger circles outlined in black indicate cluster centroids.
4.3.1. Testing the generalizability of individual differences in IBL attribute weights
The previous results showed that individual mothers are associated with different weight profiles of the attributes of the IBL instance. However, it remains unclear whether these weights are generally meaningful; that is, if they do capture individual differences that are predictive of behavioral dynamics beyond an IBL modeling context. Specifically, if these attribute weights reflect underlying individual differences in engagement and intervention dynamics, then we would expect other modeling approaches that account for these attribute weights to perform better than those that do not.
We tested this hypothesis by training and testing LSTM models with various cluster-related samples. Using the cluster assignments derived above, we randomly split each cluster in half to construct a training and a testing set. For all methods, the task is to predict the next-state engagement of the testing set during the testing period (i.e., the last 14 weeks). We examined the following methods (see  1c for a visual description):
  • Entire: One LSTM trained on all training mothers; predict for all test mothers.
  • Within-cluster: One LSTM trained on each cluster; predict for test mothers in their respective training cluster. For instance, the LSTM trained on the state-stable training set would be tested on the state-stable testing set. We hypothesize that this method should yield the most accurate predictions given the informativeness of the identified IBL clusters.
  • Outside-cluster: One LSTM trained on each cluster; predict for test mothers from a different cluster. As there are 3 clusters, we repeat this method to obtain predictions on the other non-training cluster and record the average prediction error of the two simulations. For instance, prediction accuracy for the state-stable testing set is the average of the prediction accuracy of the LSTM model trained on the transition-consistent training set and that of the model trained on the intervention-sensitive training set.
  • Random: One LSTM trained on a random subset of one-third of the training set ($N=35$); predict for all test mothers
Figure 4. Next-step prediction error per testing time step for the various LSTM training-testing methods. Note that the performance for the Outside-cluster method is an average of the performance of two LSTM models (e.g., for a state-stable mother, the average prediction error of a LSTM trained on intervention-sensitive mothers and a LSTM trained on transition-consistent mothers.
Figure 5. Comparison of prediction errors per testing mother between Entire and Within-cluster methods. Blue lines indicate mothers who are (on average) better predicted by the Within-cluster method whereas red lines indicate those better predicted by the Entire method.
Figure 4 shows the average prediction error for the various methods. In support of the generalizability of the clusters derived from IBL attribute weights, LSTMs that are trained and tested on different clusters (Outside-cluster) and a LSTM trained on a random sample (Random) perform relatively worse compared to the LSTMs that were trained and tested on the same clusters (Within).
Compared to the Entire method, the Within-cluster method does not appear to provide any additional benefit in average prediction accuracy, which might lead one to favor the simpler approach (i.e., training one vs. three models). However, as Figure 5 shows, the Within-cluster method is the more accurate one for two-thirds of the mothers (Within: 70 vs. Entire: 35).
Follow-up exploratory analyses with the Within-cluster method suggest that mothers closer to their cluster’s centroid are better predicted by the Within-cluster method as compared to the Entire method. Three binomial regression models were constructed, one for each cluster; for a given mother, the Euclidean distance between her IBL model’s attribute weights and their cluster’s centroid (dist) was used to predict the method that resulted in a lower prediction error. Although none of the effects were significant due to the imbalanced and small classes, for State-stable ($\beta_{dist}=-0.8637$) and Transition-consistent mothers ($\beta_{dist}=-1.235$), the further they were from the cluster centroid, the less likely they would be better predicted by the Within-cluster method (see Figure 6).
Figure 6. Probabilities of the Within-cluster vs. Entire method resulting in a lower prediction error. If a mother is better predicted by the Within-cluster method, she is assigned a probability of 1. Lines show predictions by binomial regression models, and shaded areas reflect the 95% CIs.
5. Conclusions and Future work
In programs with limited resources to provide interventions to a large number of beneficiaries, a key question is to decide which subset of beneficiaries should receive an intervention. In this paper, we develop a novel approach that uses computational cognitive models based on Instance-based Learning Theory (Gonzalez et al., 2003; Gonzalez and Dutt, 2011) to model the human dynamic transitions of being engaged or disengaged in a program. Given the non-Markovian nature of the transitions of a beneficiary between states (Danassis et al., 2023), a cognitive algorithm that reflects the effects of time, attention, and context similarity on human memory and decisions would simultaneously provide more accurate predictions of beneficiaries’ engagement dynamics and more meaningful individual differences that can also be leveraged by other prediction methods.
We tested our method on real-world engagement data from mothers enrolled in a maternal health program, and demonstrated that it results in higher next-step prediction accuracy compared to existing time-series methods, such as LSTM models. The personalized IBL models revealed that mothers could be clustered into three types based on their IBL attribute weights, specifically, how fluctuations in their behaviors could be attributed to intervention-sensitivity and transition-consistency. These clusters were also shown to improve the prediction accuracy of other time-series models; LSTMs trained and tested on within-cluster data outperformed all other LSTM training methods.
Future work could explore personalizing interventions and intervention schedules based on these identified individual differences in behavioral characteristics. For example, intervention-sensitive beneficiaries could be allocated more frequent but shorter calls, while transition-consistent beneficiaries could be allocated fewer but longer calls to ensure that their engagement level is maximized on each call.
One limitation of this work is the lack of an unbiased counterfactual generator to simulate the effect of recommended interventions; in our work, we used a LSTM model trained on the full dataset to generate counterfactuals. It is likely that the sequence of engagement states generated by this LSTM model does not reflect the true engagement dynamics of the actual mothers, and that the generated counterfactual states would be better predicted by another LSTM model due to the similarities in model architecture. This possibly explains the apparent contradiction between the IBL models outperforming the LSTM models on prediction accuracy (which does not involve counterfactuals) but underperforming on number of engaged mothers (which requires generating counterfactuals). Future work should prioritize testing our method with other datasets, particularly with a paradigm where the effects of recommended interventions on participants’ task engagement can be experimentally measured, rather than relying on potentially biased counterfactuals.
Another limitation is that our method currently makes predictions only from the 26th week of a mother’s pregnancy onward, as the first 25 weeks are needed to train and personalize the IBL model. A possible extension of this work could involve exploring hybrid methods, such as initially relying on predictions from other trained IBL models and gradually transitioning to the personalized model.
Acknowledgements.
This research was supported by the AI Research Institutes Program funded by the National Science Foundation under AI Institute for Societal Decision Making (AI-SDM), Award No. 2229881
Appendix A Real-world ARMMAN Dataset
A.1. Secondary Analysis
Our experiment falls into the category of secondary analysis of the data shared by ARMMAN. This paper does not involve the deployment of the proposed algorithm or any other baselines to the service call program. As noted earlier, the experiments are secondary analysis with approval from the ARMMAN ethics board.
A.2. Consent and Data Usage
Consent is obtained from every beneficiary enrolling in the NGO’s mobile health program. The data collected through the program is owned by the NGO and only the NGO is allowed to share data. In our experiments, we use anonymized call listenership logs to calculate empirical transition probabilities. No personally identifiable information (PII) is available to us. The data exchange and usage were regulated by clearly defined exchange protocols including anonymization, read-access only to researchers, restricted use of the data for research purposes only, and approval by ARMMAN’s ethics review committee.
A.3. Interventions in ARMMAN data
The interventions in ARMMAN data are chosen based on restless multi-arm bandit (RMAB) algorithms (Mate et al., 2022; Zhao et al., 2024). RMABs are a model for sequentially distributing scarce resources to a set of agents (Whittle, 1988; Xiong and Li, 2023; Jaques et al., 2019). Concretely, we have a set of arms and a limited budget and face the question of deciding which arms to pull in each round. The state of arms evolves according to a Markov Decision Process where transition probabilities depend on whether the arm is pulled in this step. RMABs have a broad range of applications, including resource allocation in anti-poaching, machine maintenance, cellular networks (Modi et al., 2019; Zhao et al., 2023). RMABs have especially be used in healthcare settings such as call scheduling in a maternal and child care program (Behari et al., 2024; Verma et al., 2024), screening patients at risk of cancer (Lee et al., 2019), and allocating hepatitis C treatment (Ayer et al., 2019).
Figure 7. Percentage of mothers engaged in the program per testing time step. $N$ = 210, budget $k$ = 6.